is x=1 a solution to 4x-3=7,yes or no

Answer by jim_thompson5910(35256)



4x-3=7


4(1)-3=7


4-3=7


1=7


Since the last equation is NOT true, this means that x = 1 is NOT a solution.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim